Sol Summary – February 08th

Crew 221 Sol Summary Report 08-Feb-2020


Summary Title: Well done…

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol Summary: Crew 221 played the rover game last night from the Planetary Volatiles Lab at York University, from Crew Journalist H. Kalucha before heading to an early night’s sleep.

Today, half the crew ( J.Todd, R.Alshammari, G. Alotaibi) went on an EVA towards Candor Chasma ,walking along Stream Bed Connector. The weather was beautiful and sunny and actually, the team reported that they were a bit too warm under the space suits. They did not achieve all their objectives but identified places to use the rover on future EVAs.

The other half of the team was staying in the hab. L. Monge and H.Kalucha were working in the science dome on soil health experiments, investigating the quality of the Martian soil being used to grow black corn. They attempted to determine if the black corn and mung beans have a chance to survive.

Commander M.Grulich was working on the oven issue, reading through the manuals and communicating with mission support to find a solution to our high propane measurements in the upper deck. The hypothesis was that the oven was not configured to be used with propane gas as the gas pressure regulator showed no change in configuration. Mission support was very helpful, checking information about the oven speedily, and eventually came to the same hypothesis.

In the afternoon, Crew 221 got external support and together with Dan, configured the oven to be used for propane. Since then, the flames of the stove are way smaller and the oven does not trigger the gas detector anymore. There are also no visible flames in the oven anymore, so the issue is solved!

This afternoon, the team worked on planting mung beans in the Greenhab that hopefully sprout soon. The crew already set a jar of mung beans aside that are sprouting since today to be hopefully ready to eat within the next couple of days.

Crew Engineer worked on his antenna project and managed to receive some signals from Arizona. Last but not least the microbiology experiments seemed to show first signs of growth on the soil samples. Tonight we will have Costa Rican pinto with a mocktail and the Costa Rican Cultural night, which we are all very excited about!

Look Ahead Plan: Sports Session, Meditation Session, Cultural night Costa Rica, Preparation rover and further microbiology samples and ISRU experiments

Anomalies in work: None

Update mouse in hab: still no joy in catching it.

Weather: sunny no winds in the morning, windy in the afternoon, no precipitation

Crew Physical Status: Good

EVA: 10:00an to 13:00pm

Reports to be filed:

Sol Summary

Commander Report

Photo Report

EVA Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site